OpenStudy (anonymous):

Find the Surface Area of the hyperbolic paraboloid f(x,y)= xy on the domain of the disk x^2 + y^2 <= 1 using polar coordinates.

OpenStudy (anonymous):

Calc 3?

OpenStudy (anonymous):

\[ f(x,y)= xy \to r^2\cos\theta\sin\theta \\ x^2 + y^2 <= 1 \to 0\leq r\leq 1,0\leq \theta\leq 2\pi \]

OpenStudy (anonymous):

\[ \Phi(r,\theta)=(x(r,\theta),y(r,\theta),z(r,\theta)) = (r\cos\theta,r\sin\theta,r^2\cos\theta\sin\theta) \\ T_r\times T_\theta =\det\begin{pmatrix} \mathbf{i}&\mathbf{j}&\mathbf{k}\\ \cos\theta &\sin\theta &2r\cos\theta\sin\theta \\ -r\sin\theta & r\cos\theta & r^2(\cos^2\theta-\sin^2\theta) \end{pmatrix} \\ ||T_r\times T_\theta || = \sqrt{(T_r\times T_\theta) \cdot (T_r\times T_\theta) } \]\[ \iint_SdS=\iint_D||T_r\times T_\theta||drd\theta \quad D:(r,\theta)\in[0,1]\times [0,2\pi] \]

OpenStudy (anonymous):

The problem is I keep getting zero. \[\cos ^{2} 2\pi - \cos ^{2} 0 = 0\]

OpenStudy (anonymous):

I'd need to see more of your work.